# Robust Nonlinear Least Squares Problems

### v1

@@ 2 The value function
which is an $m\times m$ matrix. If we project, since ${\bf P}$ is an orthogonal projector and ${\bf P}{\bf C}={\bf C}$ we have
@@
$\begin{array}[]{ll}\|{\bf P}{\bf F}({\bf x})-{\bf C}{\bf y}\|^{2}&=\|{{\bf Q}}{{\bf Q}}^{\top}{\bf F}({\bf x})-{{\bf Q}}{{\bf R}}{\bf y}\|^{2}\\ \\ &=\|{{\bf Q}}^{\top}{\bf F}({\bf x})-{{\bf R}}{\bf y}\|^{2}.\end{array}$
@@
We can solve this problem explicitly. Since ${\bf R}$ is diagonal with positive diagonal elements $R_{ii}$, we obtain
@@
The last equality is a consequence of the diagonality of ${\bf R}$ and the fact that ${\bf R}{\bf Q}^{\top}={\bf C}^{\top}$. The term $\|{\bf C}^{\top}{\bf F}({\bf x})\|_{1}$ is nonsmooth, even if ${\bf F}$ is differentiable.
We summarize the results.
Theorem 2.1.
From the expression (5), the function $\varphi$ has the following properties.
If $\mathbf{F}$ is linear, then $\varphi$ is convex.
@@
$\varphi(\mathbf{x};\mathbf{d})=\max_{\mathbf{y}\in\Omega^{*}(\mathbf{x})}\mathbf{d}^{\top}{\bf F}^{\prime}(\mathbf{x})^{\top}(\mathbf{F}(\mathbf{x})-\mathbf{y}),$
@@
If $\|\mathbf{F}(\mathbf{x})\|^{2}$ has a minimizer $\mathbf{x}^{*}$ and there is $\gamma>0$ such that for any $\mathbf{x}\not\in{\bf X}_{\gamma}:=\{\mathbf{x}\in\mathbb{R}^{n}:\,\|\mathbf{x}-\mathbf{x}^{*}\|\leq\gamma\}$, $\|{\bf C}\mathbf{F}(\mathbf{x})\|_{1}>\|{\bf C}\mathbf{F}(\mathbf{x}^{*})\|_{1}$, then $\varphi$ has a minimizer in ${\bf X}_{\gamma}$, which can be seen from the continuity of $\varphi$ and boundedness of ${\bf X}_{\gamma}$ with
@@
$\geq$ $2\delta(\|{\bf C}\mathbf{F}(\mathbf{x})\|_{1}-\|{\bf C}\mathbf{F}(\mathbf{x}^{*})\|_{1})>0,\quad\mathbf{x}\not\in{\bf X}_{\gamma}$
@@
Let Assumption 2.1 hold. If $\varphi$ has a minimizer, then problem (4) has a minimax point and $(\mathbf{x}^{*},\mathbf{y}^{*})$ is a minimax point of (4) for any
@@
Now we show $(\mathbf{x}^{*},\mathbf{y}^{*})$ is a minimax point of (4). From (19), we have $\max_{\mathbf{y}\in\Omega}f(\mathbf{x}^{*},\mathbf{y})\leq f(\mathbf{x}^{*},\mathbf{y}^{*}).$ On the other hand, we have
@@
Hence $(\mathbf{x}^{*},\mathbf{y}^{*})$ is a minimax point of (4).
From Theorem 2.2, we can develop algorithms for finding a minimax point $(\mathbf{x}^{*},\mathbf{y}^{*})$ of (4) by the following two steps:
Step 1
Find $\mathbf{x}^{*}\in\mathop{\mathrm{argmin}}_{\bf x}\varphi(\mathbf{x}).$
Step 2
Find $\mathbf{y}^{*}\in\mathop{\mathrm{argmax}}_{\mathbf{y}\in\Omega}f(\mathbf{x}^{*},\mathbf{y}).$
3 Complexity for minimax problem (4)
In this section, we show that the methods from [9] can be applied to find a first order $\epsilon$-approximate necessary minimax point of problem (4) with an optimal $O(|\log\epsilon|+\epsilon^{-2})$ evaluation bound. We assume that ${\bf F}$ is Lipschitz continuously differentiable in this section.
We consider the following minimization problem
$\min_{\bf x}\psi({\bf x}):=\|{\bf F}({\bf x})\|^{2}+2\delta\|{\bf C}^{\top}{\bf F}({\bf x})\|_{1}.$ (20)
Let the linearization of $\psi$ at $\mathbf{x}$ be defined as follows
From [9], one can verify that if ${\bf x}$ is a minimizer of problem (20) then
$\min_{\|{\bf s}\|\leq 1}\phi({\bf x},{\bf s})=\phi({\bf x},0).$ (21)
Following the definition in [9], we say that ${\bf x}$ is a first order $\epsilon$-approximate necessary minimizer of problem (20) if
$\phi({\bf x},0)-\min_{\|{\bf s}\|\leq 1}\phi({\bf x},{\bf s})=\max_{\|{\bf s}\|\leq 1}\Delta\phi({\bf x},{\bf s})\leq\epsilon,$ (22)
with
$\Delta\phi({\bf x},{\bf s})$ $=$ $\phi({\bf x},0)-\phi({\bf x},{\bf s})$
$=$ $-2({\bf F}^{\prime}({\bf x})^{\top}{\bf F}({\bf x}))^{\top}{\bf s}+2\delta(\|{\bf C}^{\top}{\bf F}({\bf x})\|_{1}-\|{\bf C}^{\top}({\bf F}({\bf x})+{\bf F}^{\prime}({\bf x}){\bf s})\|_{1}).$
We say that ${\bf x}$ is a first order necessary minimizer of problem (20) if $\epsilon=0$ in (22).
In [9], Gratton et al proposed an adaptive regularization algorithm for composite nonsmooth nonconvex optimization problems. Problem (20) satisfies Assumption AS.1-AS.4 in [9] when $\mathbf{F}$ is continuously differentiable and the level set $\{\mathbf{x}\,:\,\psi(\mathbf{x})\leq\psi(\mathbf{x}_{0})\}$ is bounded, where $\mathbf{x}_{0}$ is the initial point of the algorithm. Hence, we can use the algorithm to find a first order $\epsilon$-approximate necessary minimizer ${\bf x}$ satisfying (22), which needs at most $O(|\log\epsilon|+\epsilon^{-2})$ evaluations of ${\bf F}({\bf x})$ and ${\bf F}^{\prime}({\bf x})$.
Definition 3.1.
We say $(\hat{{\bf x}},\hat{{\bf y}})$ is a first order $\epsilon$-approximate necessary minimax point of the min-max problem (4) if
$\min_{\|{\bf s}\|\leq 1}\Delta\phi(\hat{{\bf x}},{\bf s})\leq\epsilon,$ (23)
$\hat{\mathbf{y}}\in\mathop{\mathrm{argmax}}_{{\bf y}\in\Omega}\|{\bf F}(\hat{{\bf x}})-{\bf C}^{\top}{\bf y}\|^{2}.$ (24)
We say $(\hat{{\bf x}},\hat{{\bf y}})$ is a first order necessary minimax point of the minimax problem (4) if $\epsilon=0$ in (23).
Using the results in § 2, we can solve the maximization problem (24) explicitly. Let ${\bf C}={\bf Q}{\bf R}$. Then
$\hat{\mathbf{y}}_{i}=\delta\left\{\begin{array}[]{ll}-1&\mbox{if $({\bf Q}^{\top}{\bf F}(\hat{\mathbf{x}}))_{i}>0$}\\ \{-1,1\}&\mbox{if $({\bf Q}^{\top}{\bf F}(\hat{\mathbf{x}}))_{i}=0$}\\ 1&\mbox{if $({\bf Q}^{\top}{\bf F}(\hat{\mathbf{x}}))_{i}<0$,}\end{array}\right.\quad i=1,\ldots,r.$
Hence, by Theorem 2.2 and the complexity results for minimization problem [9], we obtain that finding a first order $\epsilon$-approximate necessary minimax point of (4) needs at most $O(|\log\epsilon|+\epsilon^{-2})$ evaluations of ${\bf F}({\bf x})$ and ${\bf F}^{\prime}({\bf x})$.
4 Consequences of (5)
In this section, we use (5) to establish some interesting relations between the nonlinear least squares problem (1) and the robust nonlinear least squares problem (4). Denote the solution set of (1) and the set of minimizers of $\varphi$ defined in (5), respectively, by
@@ 4.1 $L^{1}$ optimality
$\begin{array}[]{ll}\varphi({\hat{\bf x}})&=\|{\bf F}({\hat{\bf x}})\|^{2}+2\delta\|{\bf C}^{\top}{\bf F}({\hat{\bf x}})\|_{1}+\|{\bf C}\|_{F}^{2}\delta^{2}\\ &\leq\varphi({{\bf x}^{*}})=\|{\bf F}({{\bf x}^{*}})\|^{2}+2\delta\|{\bf C}^{\top}{\bf F}({{\bf x}^{*}})\|_{1}+\|{\bf C}\|_{F}^{2}\delta^{2}\\ &\leq\|{\bf F}({\hat{\bf x}})\|^{2}+2\delta\|{\bf C}^{\top}{\bf F}({{\bf x}^{*}})\|_{1}+\|{\bf C}\|_{F}^{2}\delta^{2}.\end{array}$ (25)
@@
If ${\bf F}({\bf x}^{*})=0$ then the nonsmooth term
$02\delta\|{\bf C}^{\top}{\bf F}({\bf x}^{*})\|_{1}=$ 0 .
Hence the two solution sets ${\bf X}^{*}$ and $\hat{{\bf X}}$ are the same. From
$\mathbf{y}^{*}\in\mathop{\mathrm{argmin}}_{{\bf y}\in\Omega}\|{\bf F}({\bf x}^{*})-{\bf C}^{\top}{\bf y}\|^{2}=\mathop{\mathrm{argmin}}_{{\bf y}\in\Omega}\|{\bf C}^{\top}{\bf y}\|^{2},$
we have ${\bf y}^{*}=\delta\mathbf{c}$, where $\mathbf{c}$ is a vector with all elements being 1 or -1 under Assumption 2.1. Summarizing, we have the following theorem.
Theorem 4.1.
Let Assumption 2.1 hold. If ${\bf F}({\bf x}^{*})=0$, then $({\bf x}^{*},\delta\mathbf{c})$ is a minimax point of (4), where $\mathbf{c}$ is a vector with all elements being 1 or -1.
However, as we saw in the example in subsection 1.1, $({\bf x}^{*},\delta\mathbf{c})$ need not be a saddle point, a local saddle point, a local minimax point or a stationary point of (4). Moreover, solutions of the first order conditions (7) or (10) may be not a minimax point of (4).
@@ 4.2 Error bounds: linear least squares problems
We first consider linear least squares problem,
@@
Since the objective function of problem (27) is piecewise quadratic convex with nonnegative function values, problem (27) has a minimizer [7].
@@
$\subseteq$ $[-\|{\bf C}^{\top}{\bf A}\|_{\infty}{\bf e},\|{\bf C}^{\top}{\bf A}\|_{\infty}{\bf e}].$
where ${\bf B}_{i}$ is the $i$th row of ${\bf B}$.
@@
${\rm dist}(\mathbf{x}^{*},\hat{{\bf X}})\leq 2\delta\gamma^{2}\sqrt{r}\|{\bf C}^{\top}{\bf A}\|_{\infty}.$ (28)
Proof 4.3.
From Theorem 2.7 in [12], we have
${\rm dist}(\mathbf{x}^{*},\hat{{\bf X}})\leq\gamma\sqrt{\varphi(\mathbf{x}^{*})-\varphi(\bar{{\bf x}})},$ (29)
if $\varphi(\mathbf{x}^{*})-\varphi_{\rm min}\leq\Gamma$.
Let $\bar{\mathbf{x}}\in\hat{{\bf X}}$ such that $\|\mathbf{x}^{*}-\bar{\mathbf{x}}\|={\rm dist}(\mathbf{x}^{*},\hat{{\bf X}}).$ Since $\varphi$ is convex, we have
@@
$\subseteq$ $2\delta[-\|{\bf C}^{\top}{\bf A}\|_{\infty}{\bf e},\|{\bf C}^{\top}{\bf A}\|_{\infty}{\bf e}].$
@@
${\rm dist}(\mathbf{x}^{*},\hat{{\bf X}})\leq\gamma^{2}\|g(\mathbf{x}^{*})\|\leq 2\delta\gamma^{2}\sqrt{r}\|{\bf C}^{\top}{\bf A}\|_{\infty}.$
@@
${\rm dist}(\mathbf{x}^{*},\hat{{\bf X}})\leq\|\mathbf{x}^{*}-\hat{\mathbf{x}}\|\leq 2\delta r\|({\bf A}^{\top}{\bf A})^{\dagger}({\bf C}^{\top}{\bf A})^{\top}\|_{1}$
in a neighborhood of the minimum norm solution $\mathbf{x}^{*}$ of the linear least squares problem when $\varphi$ is differentiable at $\mathbf{x}^{*}$. This special case is different from the one in Equation (24) (page 1044) of [8]. Note that in [8], the set $\Omega=\{\mathbf{y}\,:\,\|\mathbf{y}\|\leq\delta\}$ is defined by the Euclidean norm and a linear least squares solution is a robust linear least squares solution.
@@ 4.3 Error bounds: nonlinear least squares problems
We use the second order growth condition in [1, Section 4.4.1] to present an error bound from a nonlinear least squares solution to the robust nonlinear least squares solution set.
We say that the second order growth condition holds at $\hat{{\bf X}}$ if there exists a neighborhood ${\cal N}$ of $\hat{{\bf X}}$ and a constant $\mu>0$ such that
$\varphi(\mathbf{x})\geq\varphi_{\rm min}+\mu[{\rm dist}(\mathbf{x},\hat{{\bf X}})]^{2},\,\,\,\forall\mathbf{x}\,\in{\cal N},$ (34)
where $\varphi_{\rm min}=\min_{\mathbf{x}}\varphi(\mathbf{x}).$ If $\varphi$ is a piecewise convex quadratic function, then (34) holds [12].
We say that ${\bf C}\mathbf{F}$ is Lipschitz continuous in ${\cal N}$ at $\hat{{\bf X}}$ if there is $L>0$ such that
$\|{\bf C}(\mathbf{F}(\mathbf{x})-\mathbf{F}(\mathbf{z}))\|\leq L\|\mathbf{x}-\mathbf{z}\|,\quad\quad\forall\mathbf{x}\,\in{\cal N},\mathbf{z}\in\hat{{\bf X}}.$ (35)
@@
${\rm dist}(\mathbf{x}^{*},\hat{{\bf X}})\leq 2\delta\sqrt{r}L\mu^{-1}.$ (36)
Proof 4.8.

### v2

@@ 2 The value function
which is an $m\times m$ matrix. Since ${\bf P}$ is an orthogonal projector and ${\bf P}{\bf C}={\bf C}$, we have
@@
$\begin{array}[]{ll}\|{\bf P}{\bf F}({\bf x})-{\bf C}{\bf y}\|^{2}&=\|{{\bf Q}}{{\bf Q}}^{\top}{\bf F}({\bf x})-{{\bf Q}}{{\bf R}}{\bf y}\|^{2}\\ &=\|{{\bf Q}}^{\top}{\bf F}({\bf x})-{{\bf R}}{\bf y}\|^{2}.\end{array}$
@@
We can solve this problem explicitly. Since ${\bf R}$ is diagonal with positive diagonal elements ${\bf R}_{ii}$, we obtain
@@
The last equality is a consequence of the diagonality of ${\bf R}$ and the fact that ${\bf R}{\bf Q}^{\top}={\bf C}^{\top}$.
It is worth noting that the inner function $\varphi$ is nonsmooth due to the term $\|{\bf C}^{\top}{\bf F}({\bf x})\|_{1}$, even if ${\bf F}$ is differentiable.
Remark 2.1 From the expression (5), the function $\varphi$ has the following properties.
@@
$\varphi(\mathbf{x};{\bf{d}})=\max_{\mathbf{y}\in\Omega^{*}(\mathbf{x})}{\bf{d}}^{\top}{\bf F}^{\prime}(\mathbf{x})^{\top}(\mathbf{F}(\mathbf{x})-\mathbf{y}),$
@@
If $\|{\bf C}^{\top}\mathbf{F}(\mathbf{x})\|_{1}$ has a minimizer $\mathbf{x}^{*}$ and there is $\gamma>0$ such that for any $\mathbf{x}\not\in{\bf X}_{\gamma}:=\{\mathbf{x}\in\mathbb{R}^{n}:\,\|\mathbf{x}-\mathbf{x}^{*}\|\leq\gamma\}$, $\|\mathbf{F}(\mathbf{x})\|^{2}>\|\mathbf{F}(\mathbf{x}^{*})\|^{2}$, then $\varphi$ has a minimizer in ${\bf X}_{\gamma}$, which can be seen from the continuity of $\varphi$ and boundedness of ${\bf X}_{\gamma}$ with
@@
$\geq$ $\|\mathbf{F}(\mathbf{x})\|^{2}-\|\mathbf{F}(\mathbf{x}^{*})\|^{2}>0,\quad\mathbf{x}\not\in{\bf X}_{\gamma}$
@@
The function $\|{\bf C}^{\top}\mathbf{F}(\mathbf{x})\|_{1}$ can have infinitely many minimizers, but only the minimizers at which $\|\mathbf{F}(\mathbf{x})\|$ has small values are interesting for the robust least squares problems. For example, $\mathbf{F}(\mathbf{x})=(\mathbf{x},1-\mathbf{x})^{\top}$ for ${\bf x}\in\mathbb{R}$ and ${\bf C}\in\mathbb{R}^{2\times 2}$ is the identity matrix. Any point $\mathbf{x}^{*}\in[0,1]$ is a minimizer of $\|\mathbf{F}(x)\|_{1}$, but only at $\mathbf{x}^{*}=\frac{1}{2}$, $\|\mathbf{F}(\mathbf{x}^{*})\|<\|\mathbf{F}(\mathbf{x})\|$ for $\mathbf{x}\not\in{\bf X}_{\gamma}=\{\mathbf{x}^{*}\}$.
If $\mathbf{F}$ is linear, then $\varphi$ is convex and $\mathbf{x}^{*}$ is a minimizer of $\varphi$ if and only if
$0\in\nabla\|\mathbf{F}(\mathbf{x}^{*})\|^{2}+2\delta\partial\|{\bf C}^{\top}F(\mathbf{x}^{*})\|_{1}.$ (19)
[…]
@@
Let Assumption 2.1 hold. If $\varphi$ has a minimizer, then problem (2) has a global minimax point and $(\mathbf{x}^{*},\mathbf{y}^{*})$ is a global minimax point of (2) for any
@@
Now we show $(\mathbf{x}^{*},\mathbf{y}^{*})$ is a global minimax point of (2). From (20), we have $\max_{\mathbf{y}\in\Omega}f(\mathbf{x}^{*},\mathbf{y})\leq f(\mathbf{x}^{*},\mathbf{y}^{*}).$ On the other hand, we have
@@
Hence $(\mathbf{x}^{*},\mathbf{y}^{*})$ is a global minimax point of (2).
3 Solution sets of problems (1) and (2)
In this section, we study some interesting relations between the nonlinear least squares problem (1) and the robust nonlinear least squares problem (2) by using (5). Denote the solution set of (1) and the set of minimizers of $\varphi$ defined in (5), respectively, by
@@ 3.1 Zero residual
$\begin{array}[]{ll}\varphi({\hat{\bf x}})&=\|{\bf F}({\hat{\bf x}})\|^{2}+2\delta\|{\bf C}^{\top}{\bf F}({\hat{\bf x}})\|_{1}+\|{\bf C}\|_{F}^{2}\delta^{2}\\ \\ &\leq\varphi({{\bf x}^{*}})=\|{\bf F}({{\bf x}^{*}})\|^{2}+2\delta\|{\bf C}^{\top}{\bf F}({{\bf x}^{*}})\|_{1}+\|{\bf C}\|_{F}^{2}\delta^{2}\\ \\ &\leq\|{\bf F}({\hat{\bf x}})\|^{2}+2\delta\|{\bf C}^{\top}{\bf F}({{\bf x}^{*}})\|_{1}+\|{\bf C}\|_{F}^{2}\delta^{2}.\end{array}$ (21)
@@
Proposition 3.1.
Let Assumption 2.1 hold. If ${\bf F}({\bf x}^{*})=0$, then $({\bf x}^{*},\delta{\bf{c}})$ is a global minimax point of (2), where ${\bf{c}}$ is a vector with all elements being 1 or -1.
Proof 3.2.
If ${\bf F}({\bf x}^{*})=0$ then the nonsmooth term $02\delta\|{\bf C}^{\top}{\bf F}({\bf x}^{*})\|_{1}=$ 0 . Hence the two solution sets ${\bf X}^{*}$ and $\hat{{\bf X}}$ are the same from (21). Moreover, from
$\mathbf{y}^{*}\in\mathop{\mathrm{argmax}}_{{\bf y}\in\Omega}\|{\bf F}({\bf x}^{*})-{\bf C}^{\top}{\bf y}\|^{2}=\mathop{\mathrm{argmax}}_{{\bf y}\in\Omega}\|{\bf C}^{\top}{\bf y}\|^{2},$
we have ${\bf y}^{*}=\delta{\bf{c}}$, where ${\bf{c}}$ is a vector with all elements being 1 or -1.
However, as we saw in the example in subsection 1.1, $({\bf x}^{*},\delta{\bf{c}})$ may not be a saddle point, a local saddle point, a local minimax point or a stationary point of (2). Moreover, solutions of the first order conditions (7) or (10) may not be global minimax points of (2).
@@ 3.2 Error bounds: linear least squares problems
We first consider the linear least squares problem with
@@
Since the objective function of problem (23) is piecewise quadratic convex with nonnegative function values, problem (23) has a minimizer (Frank & Wolfe, 1956).
@@
$\subseteq$ $[-|{\bf B}|^{\top}{\bf e},|{\bf B}|^{\top}{\bf e}],$
where ${\bf B}_{i}$ is the $i$th row of ${\bf B}$, $|{\bf B}|=(|{\bf B}_{i,j}|)\in\mathbb{R}^{r\times n}$, and ${\bf e}\in\mathbb{R}^{r}$ is the vector with all elements being one.
@@
${\rm dist}(\mathbf{x}^{*},\hat{{\bf X}})\leq 2\delta\gamma^{2}\sqrt{r}\|{\bf C}^{\top}{\bf A}\|.$ (24)
Proof 3.4.
From Theorem 2.7 in Li (1995), we have
${\rm dist}(\mathbf{x}^{*},\hat{{\bf X}})\leq\gamma\sqrt{\varphi(\mathbf{x}^{*})-\varphi_{\rm min}},$ (25)
if $\varphi(\mathbf{x}^{*})-\varphi_{\rm min}\leq\Gamma$. Let $\bar{\mathbf{x}}\in\hat{{\bf X}}$ such that $\|\mathbf{x}^{*}-\bar{\mathbf{x}}\|={\rm dist}(\mathbf{x}^{*},\hat{{\bf X}}).$ Since $\varphi$ is convex, we have
@@
$\subseteq$ $2\delta[-|{\bf B}|^{\top}{\bf e},|{\bf B}|^{\top}{\bf e}].$
@@
${\rm dist}(\mathbf{x}^{*},\hat{{\bf X}})\leq\gamma^{2}\|g(\mathbf{x}^{*})\|\leq 2\delta\gamma^{2}\sqrt{r}\|{\bf C}^{\top}{\bf A}\|.$
@@
${\rm dist}(\mathbf{x}^{*},\hat{{\bf X}})\leq\|\mathbf{x}^{*}-\hat{\mathbf{x}}\|\leq 2\delta r\|({\bf A}^{\top}{\bf A})^{\dagger}({\bf C}^{\top}{\bf A})^{\top}\|_{1}$ (30)
in a neighborhood of the minimum norm solution $\mathbf{x}^{*}$ of the linear least squares problem when $\varphi$ is differentiable at $\mathbf{x}^{*}$. This special case is different from the one in Equation (24) of Ghaoui & Lebret (1997, page 1044). Note that in Ghaoui & Lebret (1997), the set $\Omega=\{\mathbf{y}\,:\,\|\mathbf{y}\|\leq\delta\}$ is defined by the Euclidean norm and a linear least squares solution is a robust linear least squares solution.
@@ 3.3 Error bounds: nonlinear least squares problems
We use the second order growth condition in Bonnans & Shapiro (2000, Section 4.4.1) to present an error bound from a nonlinear least squares solution to the robust nonlinear least squares solution set.
We say that the second order growth condition holds at $\hat{{\bf X}}$ if there exists a neighborhood ${\cal N}$ of $\hat{{\bf X}}$ and a constant $\alpha>0$ such that
$\varphi(\mathbf{x})\geq\varphi_{\rm min}+\alpha[{\rm dist}(\mathbf{x},\hat{{\bf X}})]^{2},\,\,\,\forall\mathbf{x}\,\in{\cal N},$ (31)
where $\varphi_{\rm min}=\min_{\mathbf{x}}\varphi(\mathbf{x}).$ If $\varphi$ is a piecewise convex quadratic function, then (31) holds (Li, 1995).
We say that ${\bf C}^{\top}\mathbf{F}$ is Lipschitz continuous in ${\cal N}$ at $\hat{{\bf X}}$ if there is $L>0$ such that
$\|{\bf C}^{\top}(\mathbf{F}(\mathbf{x})-\mathbf{F}(\mathbf{z}))\|\leq L\|\mathbf{x}-\mathbf{z}\|,\quad\quad\forall\mathbf{x}\,\in{\cal N},\mathbf{z}\in\hat{{\bf X}}.$ (32)
@@
${\rm dist}(\mathbf{x}^{*},\hat{{\bf X}})\leq 2\delta\sqrt{r}L\alpha^{-1}.$ (33)
Proof 3.9.